OpenStudy (anonymous):

Whats 10+9? is it 21?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Close this now, It's spam.

OpenStudy (anonymous):

And 10 + 9 = 19.

OpenStudy (anonymous):

Why couldnt u tell me that before?

OpenStudy (anonymous):

Close the question....

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

Im taking your medal away

OpenStudy (anonymous):

I don't care about medals.

OpenStudy (anonymous):

Just close your stupid spam.

OpenStudy (anonymous):

i tink u do care

OpenStudy (anonymous):

9 + 10 is 21...

OpenStudy (anonymous):

@AnswerMyQuestions Must have woke up on the wrong side of the bed.

OpenStudy (k_lynn):

@ganeshie8 would this be considered spam?

OpenStudy (anonymous):

Actually @k_lynn Some people might be confused with this theory of what the answer to 9 + 10 is...

OpenStudy (anonymous):

Duh

OpenStudy (anonymous):

Nope. @101

OpenStudy (anonymous):

@k_lynn Of course It's spam. Who won't know the answer to 10 + 9? Some idiot made up something that 10 + 9 = 21.

OpenStudy (k_lynn):

I get that @101 but what happens is I think 10+9=21 was used in a vine...so some users probably find it funny to post this.

OpenStudy (anonymous):

And we get this spam 5+ times a day.

OpenStudy (anonymous):

Get the chill pill @AnswerMyQuestions

OpenStudy (anonymous):

Indeed. They joke and ask the dumb question.

OpenStudy (anonymous):

@AnswerMyQuestions 10+9=21 u stupid

OpenStudy (anonymous):

You wish.

OpenStudy (anonymous):

fed troll is fed

OpenStudy (anonymous):

@ganeshie8 @Abhisar @TuringTest Can one of you please close this spam?

OpenStudy (anonymous):

Lol....